calculate the percent by mass of solute of the 8.60 g of salt in 95.0 g of solution


1
Expert's answer
2021-01-06T05:55:32-0500

%(m/m) = "\frac{mass of solute }{mass of solution}" × 100%


"\frac{8.60g}{95.0g}" × 100%


= 9.052 %